OpenStudy (anonymous):

Solve d=16t ^2 to isolate t alone.

OpenStudy (midhun.madhu1987):

divide both sides by 16..wht u gt?

OpenStudy (anonymous):

d/16=t^2 right?

OpenStudy (midhun.madhu1987):

yes,,, now take square root on both sides..u get t

OpenStudy (anonymous):

d/4=t ?

OpenStudy (midhun.madhu1987):

nope ... square root of d

OpenStudy (midhun.madhu1987):